# Benchmarking Mental State Representations in Language Models

### v1

@@ 3.4 Probing Experiments
ITI $52_{+0}$ $45_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+1}}$ $35_{+0}$ $64_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+5}}$ $61_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+11}}$ $46_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+9}}$ $48_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+2}}$ $59_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}$ $42_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+9}}$
CAA $85_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+33}}^{*}$ $88_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+44}}^{*}$ $66_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+31}}^{*}$ $71_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+12}}^{*}$ $69_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+19}}^{*}$ $55_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+18}}^{*}$ $75_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+29}}^{*}$ $92_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+43}}^{*}$ $59_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+26}}^{*}$
@@
ITI $84_{+0}$ $65_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+9}}$ $59_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+12}}$ $78_{+0}$ $58_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+7}}$ $47_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+9}}^{*}$ $72_{+0}$ $60_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+12}}$ $48_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+17}}$
CAA $97_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+13}}^{*}$ $94_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+38}}^{*}$ $91_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+44}}^{*}$ $80_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+2}}^{*}$ $71_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+20}}^{*}$ $54_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+16}}^{*}$ $97_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+25}}$ $94_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+46}}^{*}$ $87_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+56}}^{*}$
@@
ITI $90_{+0}$ $90_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+3}}$ $78_{+0}$ $94_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+1}}$ $55_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+3}}$ $50_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+2}}$ $77_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+4}}$ $58_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+5}}$ $37_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+5}}$
CAA $99_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+9}}^{*}$ $97_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+10}}^{*}$ $95_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+17}}^{*}$ $94_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+1}}^{*}$ $80_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+28}}^{*}$ $73_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+25}}^{*}$ $94_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+21}}$ $92_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+39}}^{*}$ $83_{{\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}+51}}^{*}$

### v3

@@ 4 Results
ITI $52_{+0}$ $45_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+1}}$ $35_{+0}$ $64_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+5}}$ $61_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+11}}$ $46_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+9}}$ $48_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+2}}$ $59_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+10}}$ $42_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+9}}$
CAA $85_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+33}}^{*}$ $88_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+44}}^{*}$ $66_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+31}}^{*}$ $71_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+12}}^{*}$ $69_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+19}}^{*}$ $55_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+18}}^{*}$ $75_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+29}}^{*}$ $92_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+43}}^{*}$ $59_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+26}}^{*}$
@@
ITI $84_{+0}$ $65_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+9}}$ $59_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+12}}$ $78_{+0}$ $58_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+7}}$ $47_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+9}}^{*}$ $72_{+0}$ $60_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+12}}$ $48_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+17}}$
CAA $97_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+13}}^{*}$ $94_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+38}}^{*}$ $91_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+44}}^{*}$ $80_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+2}}^{*}$ $71_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+20}}^{*}$ $54_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+16}}^{*}$ $97_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+25}}$ $94_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+46}}^{*}$ $87_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+56}}^{*}$
@@
ITI $90_{+0}$ $90_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+3}}$ $78_{+0}$ $94_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+1}}$ $55_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+3}}$ $50_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+2}}$ $77_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+4}}$ $58_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+5}}$ $37_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+5}}$
CAA $99_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+9}}^{*}$ $97_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+10}}^{*}$ $95_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+17}}^{*}$ $94_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+1}}^{*}$ $80_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+28}}^{*}$ $73_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+25}}^{*}$ $94_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+21}}$ $92_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+39}}^{*}$ $83_{{\color[rgb]{0.0,0.47,0.44}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.47,0.44}+51}}^{*}$